

********************************************************************
UARS WEEKLY STATUS REPORT - 14 MAY 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was turned off for May 6-13.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1999039. MEPS data files between
1999040 and 1999079 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 01 May 2004 (Orbit 69150, GMT Day 122) through
07 May 2004 (Orbit 69254, GMT Day 128).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 123/2246 69179 POWER OFF
PEM 125/2202 69208 ZEPS/NEPS OFF
PEM 125/2202 69208 AXIS 1 & 2 OFF
SOLSTICE 126/2238 69224 SOLSTICE TO DAY ONLY
SUSIM 126/2238 69224 POWER OFF
A Forward-to-Reverse Yaw Around maneuver (number 129) was successfully
executed in Orbit 69249 beginning 128/14:51:39Z and completing at
128/15:28:49Z on 07 May 2004. All instruments were in their off or
standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 57.89 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 24.96 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 21.55 and 21.32 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 06
June 2004 (DOY 158).
The beta angle ranged between the angles of 24.1 degrees to 01.2
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 36 degrees on 22 May 2004 (DOY 143).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.29 degrees C. Additional
battery performance data for each battery on 30 April and 07 May 2004
are provided in Appendix D.
The clock error ranged from 10.5 to -12.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 01
PEM = 02
SOLSTICE = 01
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT began taking SSA only events on TDW on 23 April 2004. This
change to operations was needed due to the TDRS-8 move to TDW. Since
the SSA service is in higher demand than the MA service, the FOT
receives approximately 50% of the events requested. This effectively
yields the realtime support of 2.5 TDRS instead of 3 TDRS. The FOT
will continue to operate in this mode until the software fix to the
CMS is made to ingest the SMA view periods correctly. Preliminary
estimates project this fix to be released to operations no sooner than
two months from now.
MEETINGS
The FOT attended the UARS Monthly Status Review on 03 May 2004 to
review the significant events for March and April.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 06 June 2004 (DOY 158). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 07 May 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
69160 122/1634 04-127 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
69179 123/2246 04-128 HALOE POWER OFF
69181 124/0209 04-129 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
69208 125/2202 04-130 PEM ZEPS/NEPS OFF
69208 125/2202 04-131 PEM AXIS 1 & 2 OFF
69224 126/2238 04-132 SOLSTICE SOLSTICE TO DAY ONLY
69224 126/2238 04-133 SUSIM POWER OFF
69234 127/1438 04-134 MPS VT LEVEL 6 (YAW)
69241 128/0236 04-135 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
69248 128/1355 04-134 OBC/TAM YAW MANEUVER COMMANDS
69249 128/1525 04-134 OBC YAW MANUEVER
69249 128/1550 04-134 OBC/TAM YAW MANEUVER COMMANDS
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 07 May 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 07 May 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
69160 122/1636 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
69181 124/0209 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
69241 128/0236 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 07 May 2004
30 April 2004 - GMT Day 121 Beta = 26.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-179.2 +5.46/+4.50 +28.4/-14.0 24.6 1.04 20.4
3 +22.4/-134.4 +4.50/+3.55 +27.6/-14.4 24.6 1.03 20.2
07 May 2004 - GMT Day 128 Beta = 01.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-184.8 +5.79/+4.50 +27.2/-13.6 25.0 1.00 21.6
3 +22.4/-072.8 +4.50/+3.23 +26.8/-14.4 25.0 1.00 21.3
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
69234 127/1438 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 07 May 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-024 FOT 69223 126/2132 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-019 FOT See List Apr 04 30832 Generic Late Acquisition
04-021 FOT 69060 116/0050 30824 Telemetry Dropouts
04-022 FOT 69132 120/2000 35744 Negative Acquisition
04-023 FOT 69134 120/2331 35746 Data Drop Out
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-024 FOT See List May 04 30832 Generic Late Acquisition